OpenStudy (anonymous):

Help.?! Find the number half way between 1/2 and 5/6.

OpenStudy (mathstudent55):

The number half way between two numbers is the average of the numbers. Add the two numbers together and divide the sum by 2.

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{2} \left(\frac{5}{6}-\frac{1}{2}\right)=\frac{2}{3} \]